vincent is 10 years older than brian. Five years ago, his age was twice that of Brian. Find their present ages.
Answer by ankor@dixie-net.com(22740)

Take it step-by-step, write an expression or an equation for each statement
:
Let v = Vincent's present age and b = Brian's present age
:
"vincent is 10 years older than brian."
v = b + 10
:
"Five years ago," from this,we can write the ages as:
(v-5)
and
(b-5)
:
" Five years ago, his age was twice that of Brian. "
(v-5) = 2(b-5)
Simplify, do some basic algebra on this equation
v - 5 = 2b - 10
v = 2b - 10 + 5
v = 2b - 5
:
"Find their present ages."
From the 1st statement(v = (b+10)); substitute (b+10) for v in the above equation:
b + 10 = 2b - 5
b - 2b = -5 - 10
-b = -15
b = +15 is Brian's age now, then v's age now = 25, obviously
:
:
Check our solution in the statement:
"Five years ago, his age was twice that of Brian."
25 - 5 = 2(15-5)
:
Could you follow the procedure here? Did it make sense to you?